40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2359     Entry time: Wed Jun 5 14:16:07 2019
Author: Koji 
Type: Update 
Category: 2micronLasers 
Subject: 2um laser / cryo stat: setup inspection and action item updates 

[Aidan, Chris, Koji]

We went down to the lab to check the situation of the setups for 2um laser measurement and stabilization and the new cryostat.

[2um laser frequency noise measurement]

  • Looked at the add-on transimpedance amps: Something was wrong with them. The power bypass caps are attached to the "hot" supply lines in parallel (both sides of the caps are soldered to a same line). And some power supply lines have no voltage. This circuit is not necessary to be bipolar. To be fixed (KA)
  • We temporarily connected one of the thorlabs 2um InGaAs biased detectors to an SR560. It showed reasonable output: DC/AC response OK & no nonsense.
  • The AOM was bypassed and the homodyne fringe was checked.
    The fringe visibility was low (~10%) and was dependent on the stress applied to the delayline fiber.
    Suspected polarization rotation somewhere -> ToDo: Check the polarization states of the output beams.
  • ToDo: Check should be done with each component. how much are the output power, output polarization, dependence/fluctuation of the polarization, etc. 
    We might be able to use the 2um Faraday Isolators (as PBSs) for the measuement.
  • Checked the fringing of the fiber delayline Mach Zehnder. We observed one fringe per sec level fluctuation.
  • Laser current actuation was checked and it turned out that it is so strong and sufficient to lock the delayline fringe.

[2um AOM]

  • The fiber coupled AOM gave us a reasonable amount of DC/AC actuation of the laser intensity.
  • The power of the 1st order output has the dependence on the "freq input" of the driver. This is probably because of the matching between the fiber coupling and the deflection angle, which is freq dependent.
  • When the freq input is 8.8V_DC, the 1st order output has the maximum efficiency. The efficiency was 96%@990mV_DC input to the modulation in.
  • The AOM actuation bandwidth was tested to be ~MHz, at least.
  • We are not supposed to give more than 1V to the modulation in while we want to apply 8.8V to the freq input. Incorrect plugging may cause the damage of the modulation input port. The setup needs to be improved with a protection circuits / AOM driver circuit.
  • Our understanding is that the modulation input has a 50Ohm input impedance while the freq input has high-Z
  • The next step towards the intensity stabilization is low noise photodetector circuits and proper interface to the AOM driver.
  • Also we want to set up TECs and other circuits for the LaserComponents PDs.

[Cryostat]

  • Cleaning: there are many components are scattered on the table.
  • Plan:
    • Move the Zack rack to the next of the optical table (or somewhere)
    • Move the yellow chemical cabinet to the place where the rack was. We can pile up some plastic boxes on it.
    • Remove the delicate optics from the steel table.
    • Place heavy cryostat components on the steel table.
    • Connect the cryo cooler to the cryostat. How do we do that? Fisrt rigidly attach for testing and then move to soft attaching?
    • Replace the optical windows to the 2um ones (2"). The current ones are for 1.5um.
    • We need a 2" 50/50 BS at 2um. Lenses and steering mirrors are in hand.
ELOG V3.1.3-